# UniGraph: Learning a Cross-Domain Graph Foundation Model From Natural Language

### v1

@@ Abstract
Foundation models like ChatGPT and GPT-4 have revolutionized artificial intelligence, exhibiting remarkable abilities to generalize across a wide array of tasks and applications beyond their initial training objectives. However, when this concept is applied to graph learning, a stark contrast emerges. Graph learning has predominantly focused on single-graph models, tailored to specific tasks or datasets, lacking the ability to transfer learned knowledge to different domains. This limitation stems from the inherent complexity and diversity of graph structures, along with the different feature and label spaces specific to graph data. In this paper, we present our UniGraph framework, designed to train a graph foundation model capable of generalizing to unseen graphs and tasks across diverse domains. Unlike single-graph models that use pre-computed node features of varying dimensions as input, our approach leverages Text-Attributed Graphs (TAGs) for unifying node representations. We propose a cascaded architecture of Language Models (LMs) and Graph Neural Networks (GNNs) as backbone networks with a self-supervised training objective based on Masked Graph Modeling (MGM). We introduce graph instruction tuning using Large Language Models (LLMs) to enable zero-shot prediction ability. Our comprehensive experiments across various graph learning tasks and domains demonstrate the model’s effectiveness in self-supervised representation learning on unseen graphs, few-shot in-context transfer, and zero-shot transfer, even surpassing or matching the performance of GNNs that have undergone supervised training on target datasets.
Foundation Models, Graph Representation Learning, Pre-Training
1 Introduction
@@
Foundation models in artificial intelligence are large-scale pre-trained models that provide a versatile base, enabling a wide range of applications and tasks (Bommasani et al., 2021). However, in the graph learning community, the long-standing practice is to train a model specific to one graph at a time (Kipf & Welling, 2016a), which we call a single-graph model. These models often can only handle one or a few tasks and lack the ability to transfer to other graphs. Also, single-graph models typically require a substantial amount of labeled data for each specific task, which can be a significant limitation in data-scarce scenarios.
Challenges. Despite many successful and mature techniques regarding foundation models in the computer vision (CV) (Radford et al., 2021) and natural language processing (NLP) (Devlin et al., 2019; Brown et al., 2020) fields, applying them to the field of graph learning poses significant challenges, due to the unique characteristics of graph data. The major challenge is the diversity of graph domains.
@@
Firstly, graphs from different domains have distinct feature spaces and label spaces. In NLP, texts from different domains, despite having vastly different semantics, can still be encoded by the same dictionary. However, for graphs from different domains, their nodes and edges have different types and semantics, leading to incompatibility between their features, making it very difficult to unify model inputs. Furthermore, graph models like GNNs (Kipf & Welling, 2016a), use a static softmax classifier for prediction, lacking the capability for zero-shot prediction on unseen classes.
@@
Secondly, as a universal data structure applied across various domains, graphs exhibit significant structural differences. For instance, citation networks are graphs where nodes represent scholarly articles, and edges represent citations from one article to another. In knowledge graphs, nodes represent entities, and edges represent relations, which have diverse structural patterns depending on the specific relation. It is challenging for graph models to transfer learned structural knowledge across different domains.
@@
Presented work. Motivated by these challenges, in this work, we propose the UniGraph framework to design and train a graph foundation model capable of generalizing to a variety of tasks on unseen graphs across diverse domains.
@@
To align the feature spaces of different graphs, we use text as a unifying medium. Many graphs in the real world contain textual features, such as citation networks and Wikipedia networks, and are known as Text-Attributed Graphs (TAGs). Compared to pre-processed vector features, textual features provide a consistent representation, which is beneficial for the transferability of the model. In addition to natural TAGs, we also explore using text to represent features of non-natural TAGs, such as molecular graphs, making it possible for the model to generalize to a broader range of domains.
@@
Inspired by recent graph prompt learning approaches (Huang et al., 2023; Liu et al., 2023a), we design a universal template to unify different tasks by contextualizing the nodes/edges/graphs on which we make predictions. Some perspectives suggest that representative local graph structural patterns are universal and transferable across different graphs (Newman, 2018). For node/edge-level tasks on large-scale graphs, we adopt the Personalized PageRank (PPR) algorithm to sample subgraphs, thereby mitigating differences in graph structures across different domains, while also making the model scalable.
@@
We propose a cascaded architecture of LMs and GNNs and design new pre-training objectives based on Masked Graph Modeling (MGM), specifically tailored for self-supervised learning on TAGs at scale. To further endow our model with zero-shot capabilities, we concatenate pre-trained graph embeddings with natural language instructions and perform instruction tuning on LLMs, allowing us to unify the label spaces of graphs from different domains through natural language in a generative manner.
@@
  • We identify the challenges in developing a cross-domain graph foundation model and present the use of TAGs to unify graphs from diverse domains. We propose a novel cascaded LM and GNN architecture for large-scale self-supervised pre-training on TAGs.
@@
  • We conduct experiments on node/edge/graph-level tasks across 11 different graph datasets from 5 distinct domains, with the largest graph comprising 111 million nodes. UniGraph offers superior or comparable performance over graph-specific models.
2 Related Work
Cross-graph learning. For graph transfer learning within the same domain, the most successful are pre-trained models on molecular graphs (Liu et al., 2021), benefiting from similar node/edge semantics and structural patterns. In addition, techniques such as fine-tuning (Hou et al., 2022), domain adaptation (Dai et al., 2022), and prompt graphs (Huang et al., 2023) are also used to achieve cross-graph transfer within the same domain. Most of these methods require the alignment of different graphs in both feature space and label space. For cross-domain graph learning, GCC (Qiu et al., 2020), by ignoring node features and utilizing graph contrastive learning for pretraining structural representations, achieves transfer between academic graphs and social graphs. OFA (Liu et al., 2023a) leverages pre-trained language models to align the feature spaces of graphs from different domains, while also transforming all downstream classification tasks into binary classification tasks, enabling it to conduct supervised training across all graphs.
Figure 1: Overview of UniGraph framework. 1) In pre-training, we employ a self-supervised approach, leveraging TAGs to unify diverse graph data. This phase involves a cascaded architecture combining LMs and GNNs. We propose Graph Siamese Masked Autoencoders as the training architecture, which learns to reconstruct the masked text of each node using the text of its neighbors. 2) In few-shot transfer, the pre-trained model can make predictions with minimal data by comparing the embeddings of the query and support graphs. 3) Zero-shot transfer is achieved through graph instruction tuning with LLMs, enabling it to understand category labels in natural language and make predictions on unseen graphs without any graph-specific training.
3 Preliminaries
3.1 Text-Attributed Graphs (TAGs)
@@
A Text-Attributed Graph (TAG) is defined as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T}_{\mathcal{V}},\mathcal{T}_{\mathcal{E}})$, where $\mathcal{V}$ represents the set of nodes and $\mathcal{E}$ represents the set of edges. For each node $v\in\mathcal{V}$, there is an associated text $t_{v}\in\mathcal{T}_{\mathcal{V}}$ representing node-level textual information. For each edge $e_{vu}\in\mathcal{E}$ connecting nodes $v$ and $u$, there is an associated text $t_{e_{vu}}\in\mathcal{T}_{\mathcal{E}}$ representing edge-level textual information. It is possible for a TAG to have only $\mathcal{T}_{\mathcal{V}}$ (node-level text).
3.2 Problem Definition
This paper introduces a cross-domain graph foundation model, pre-trained in a self-supervised manner. Our model’s generalization ability is evaluated across domains using unseen datasets through three different machine learning problems. The pre-trained model, denoted as $f_{\theta}$, operates on a TAG and generates an embedding for each node. Formally, the model’s function can be expressed as: $f_{\theta}:\mathcal{G}\rightarrow\mathbb{R}^{|V|\times d}$, where $d$ is the dimensionality of the embedding space.
Self-supervised representation learning. The primary aim of self-supervised learning (SSL) is to produce embeddings that are useful in downstream tasks. We adopt a linear probing setting to evaluate the representation learning ability of a frozen pre-trained model $f_{\theta}$.
@@ 4 The UniGraph Approach
4.1 Unifying Graphs and Tasks from Diverse Domains
Graphs from different domains often have different applications, corresponding to different tasks. Graph learning tasks can generally be divided into node, edge, and graph-level tasks, each focusing on different parts of the graph. The key to using one model to handle any task on any graph lies in finding a universal function acting as a versatile mapping tool, adaptable to different graph learning tasks. Some recent works use prompt graphs (Liu et al., 2023a, b; Sun et al., 2023) to unify the three tasks, but here we find that our simpler alternative works well, as it doesn’t involve modifying the graph or adding learnable prompt tokens. In this paper, we utilize the concept of Anchor Node(s) and their contextual subgraph(s) to construct this universal function. The unification of node, edge, and graph-level tasks can be achieved through a general contextual subgraphs processing and Anchor Nodes embedding refinement operation, denoted by $\mathbf{h}=g(\mathcal{X},\mathcal{A})$, where $\mathbf{h}$ represents the output vector representation for a node, an edge or a graph, $\mathcal{X}$ denotes the set of contextual subgraphs, and $\mathcal{A}$ signifies the set of Anchor Node(s). Formally, we define the function $g$ with our pre-trained model $f_{\theta}$ using the format:
$\mathbf{h}=g(\mathcal{X},\mathcal{A})=R(f_{\theta}(\mathcal{X}),\mathcal{A}),$ (1)
@@ 4.1 Unifying Graphs and Tasks from Diverse Domains
Table 1: Statistics of all eleven text-attributed graph datasets.
Dataset Domain Task #Graphs Avg. #Nodes Avg. #Edges Raw Texts
Cora Citation Node 1 2,708 5,429 paper titles and abstracts
PubMed Citation Node 1 19,717 44,338 paper titles and abstracts
ogbn-Arxiv Citation Node 1 169,343 1,166,243 paper titles and abstracts
ogbn-Papers100M Citation Node 1 111,059,956 1,615,685,872 paper titles and abstracts
ogbn-Products Product Node 1 2,449,029 61,859,140 product descriptions
Wiki-CS Web link Node 1 11,701 216,123 wikipedia entry names and contents
FB15K237 Knowledge Edge 1 14,541 310,116 entity names and descriptions
WN18RR Knowledge Edge 1 40,943 93,003 entity names and descriptions
PCBA Molecule Graph 437,929 26.0 28.1 textual descriptions of atoms/bonds
HIV Molecule Graph 41,127 25.5 27.5 textual descriptions of atoms/bonds
ChEMBL Molecule Graph 365,065 25.9 55.9 textual descriptions of atoms/bonds
4.2 Graph Siamese Masked Autoencoders on TAGs
The existing graph self-supervised learning methods (Veličković et al., 2018b; Kipf & Welling, 2016b; Thakoor et al., 2021; Hou et al., 2022) adopt GNNs as the backbone networks and use pre-processed vector features as input. In this part, we propose an end-to-end self-supervised learning method on TAGs. We cascade a pre-trained language model (LM) and a GNN as the backbone network. In this paper, we adopt DeBERTa-base (He et al., 2020) as our LM and GAT (Veličković et al., 2018a) as our GNN. Inspired by successful self-supervised learning techniques such as masked modeling (Kenton & Toutanova, 2019; He et al., 2022) and siamese networks (Wu et al., 2018), we designed Graph Siamese Masked Autoencoders, enabling large-scale self-supervised pre-training on TAGs.
@@ 4.2 Graph Siamese Masked Autoencoders on TAGs
The masking process involves randomly replacing a subset of tokens in each $t_{v}$ with a [MASK] token, defined by the masking function:
@@
Encoder. Our encoder consists of an LM $f_{\text{LM}}$ and a GNN $f_{\text{GNN}}$. For each node $v\in\mathcal{V}$, the masked textual feature sequence $m_{v}$ is processed through the LM $f_{\text{LM}}$ to get the hidden representations: $\widetilde{{\bm{E}}}_{v}=f_{\text{LM}}(m_{v})$.
Then, we extract embedding of the [CLS] token $\widetilde{{\bm{e}}}_{v_{\text{cls}}}\in\mathbb{R}^{d}$ from the masked output $\widetilde{{\bm{E}}}_{v}\in\mathbb{R}^{(n_{v}+2)\times d}$. The GNN $f_{\text{GNN}}$ propagates embeddings of the [CLS] tokens across the graph. The input to $f_{\text{GNN}}$ is a matrix $\widetilde{{\bm{E}}}_{\text{cls}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ consisting of embedding of the [CLS] token for all nodes. The output of the GNN, $\widetilde{{\bm{E}}}^{\prime}_{\text{cls}}\in\mathbb{R}^{|\mathcal{V}|\times d}$, is another matrix of embeddings, representing the propagated features: $\widetilde{{\bm{E}}}^{\prime}_{\text{cls}}=f_{\text{GNN}}(\mathcal{G},\widetilde{{\bm{E}}}_{\text{cls}})$
Decoder. In decoding, we adopt the masked language modeling (MLM) (Kenton & Toutanova, 2019) as objective. The intuition behind designing such training objectives is that the model can learn to reconstruct the masked text of each node using the text of its neighbors, thereby fully exploring the graph structure while learning to understand the text. For each node $v$, the embeddings of masked textual feature $\widetilde{{\bm{E}}}_{v}$ from the masked forward pass and the GNN output of [CLS] token $\widetilde{{\bm{e}}}^{\prime}_{v_{\text{cls}}}$ are concatenated and linearly transformed:
${\bm{H}}_{v}=\text{Linear}(\widetilde{{\bm{E}}}_{v}\oplus(\widetilde{{\bm{e}}}^{\prime}_{v_{\text{cls}}}\otimes\mathbf{1}_{n_{v}+2}^{\top})),$ (3)
@@
$\mathcal{L}_{\text{mask}}=-\frac{1}{\sum_{v}\sum_{i}\mathbb{I}(v,i)}\sum_{v\in\mathcal{V}}\sum_{i=1}^{n_{v}}\mathbb{I}(v,i)\cdot\log\mathbf{P}_{v}[i,T_{i}]$ (4)
@@
Latent space regularization. Some recent works suggest that designing new pretext tasks in the latent space helps stabilize the training of masked autoencoders, further enhancing the quality of the latent space representations (Chen et al., 2023; Dong et al., 2022). In this work, we propose the use of a target network to impose regularization constraints on the latent space. The target network shares the same architecture as the encoder, comprising an LM $f^{\prime}_{\text{LM}}$ and a GNN $f^{\prime}_{\text{GNN}}$, with parameters not updated through gradients. The parameters of $f^{\prime}_{\text{LM}}(*;\delta^{\prime})$ and $f_{\text{LM}}(*;\delta)$ are shared, while the parameters of $f^{\prime}_{\text{GNN}}(*;\xi^{\prime})$ are updated through exponential moving average (EMA) of $f_{\text{GNN}}(*;\xi)$ using weight decay $\tau$ to avoid collapse in negative-free SSL frameworks (Grill et al., 2020; Caron et al., 2021):
$\delta^{\prime}\leftarrow\delta\quad\text{and}\quad\xi^{\prime}\leftarrow\tau\xi^{\prime}+(1-\tau)\xi.$ (5)
@@
$\mathcal{L}_{latent}=\frac{1}{|\mathcal{V}|}\sum_{i}^{|\mathcal{V}|}(1-\frac{\bar{{\bm{z}}}_{i}^{\top}{\bm{e}}^{\prime}_{{\text{cls}}_{i}}}{\lVert\bar{{\bm{z}}}_{i}\rVert\cdot\lVert{\bm{e}}^{\prime}_{{\text{cls}}_{i}}\rVert})$ (6)
@@
$\mathcal{L}=\mathcal{L}_{mask}+\lambda\cdot\mathcal{L}_{latent}.$ (7)
@@
Table 2: Experiment results in self-supervised representation learning. We report accuracy (%) for node/edge classification tasks and ROC-AUC score (%) for graph classification tasks. The performance of supervised methods, which are trained on the individual target dataset, is marked in gray. UniGraph and other self-supervised baselines (rows in white) are pretrained on ogbn-Papers100M, and then evaluated on the individual target dataset. NoPretrain represents a randomly initialized model in our framework without any pre-training.

### v2

# UniGraph: Learning a Unified Cross-Domain Foundation Model for Text-Attributed Graphs

@@ Abstract.
Foundation models like ChatGPT and GPT-4 have revolutionized artificial intelligence, exhibiting remarkable abilities to generalize across a wide array of tasks and applications beyond their initial training objectives. However, graph learning has predominantly focused on single-graph models, tailored to specific tasks or datasets, lacking the ability to transfer learned knowledge to different domains. This limitation stems from the inherent complexity and diversity of graph structures, along with the different feature and label spaces specific to graph data. In this paper, we recognize text as an effective unifying medium and employ Text-Attributed Graphs (TAGs) to leverage this potential. We present our UniGraph  [^1] The code is available at https://github.com/yf-he/UniGraph framework, designed to learn a foundation model for TAGs, which is capable of generalizing to unseen graphs and tasks across diverse domains. Unlike single-graph models that use pre-computed node features of varying dimensions as input, our approach leverages textual features for unifying node representations, even for graphs such as molecular graphs that do not naturally have textual features. We propose a novel cascaded architecture of Language Models (LMs) and Graph Neural Networks (GNNs) as backbone networks. Additionally, we propose the first pre-training algorithm specifically designed for large-scale self-supervised learning on TAGs, based on Masked Graph Modeling. We introduce graph instruction tuning using Large Language Models (LLMs) to enable zero-shot prediction ability. Our comprehensive experiments across various graph learning tasks and domains demonstrate the model’s effectiveness in self-supervised representation learning on unseen graphs, few-shot in-context transfer, and zero-shot transfer, even surpassing or matching the performance of GNNs that have undergone supervised training on target datasets.
Graph Neural Networks, Self-Supervised Learning, Graph Pre-Training
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Neural networksccs: Information systems Social networks
1. Introduction
@@
Foundation models in artificial intelligence are large-scale pre-trained models that provide a versatile base, enabling a wide range of applications and tasks (Bommasani et al., 2021). However, in the graph learning community, the long-standing practice is to train a model specific to one graph at a time (Kipf and Welling, 2016a), which we call a single-graph model. These models often can only handle one or a few tasks and lack the ability to transfer to other graphs. Also, single-graph models typically require a substantial amount of labeled data for each specific task, which can be a significant limitation in data-scarce scenarios.
@@
Challenges in building a graph foundation model. The success of language and vision foundation models is built upon invariances across their respective application domains, such as a unified vocabulary or raw pixels. The major challenge in learning a graph foundation model is the diversity of graph domains and, furthermore, how to learn the invariances across different domains. Firstly, graphs from different domains have distinct feature spaces and label spaces. In NLP, texts from different domains, despite having vastly different semantics, can still be encoded using the same dictionary, thereby generating transferable representations. However, for graphs from different domains, their nodes and edges have different types and semantics, leading to incompatibility between their features, making it very difficult to unify model inputs. Furthermore, graph models like GNNs (Kipf and Welling, 2016a), use a static softmax classifier for prediction, lacking the capability for zero-shot prediction on unseen classes in different label spaces.
@@
Secondly, as a universal data structure applied across various domains, graphs exhibit significant structural differences. Citation networks are directed and often acyclic, with nodes representing scholarly articles and edges representing citations. In contrast, knowledge graphs are more complex, with nodes representing entities and diverse edges representing relations, often forming intricate patterns like cycles and cliques. It is challenging for graph models to transfer learned structural knowledge across different domains.
@@
Presented work. Motivated by these challenges, in this work, we propose the UniGraph framework to design and train a foundation model capable of generalizing to a variety of tasks on unseen graphs across diverse domains.
@@
To align the feature spaces of different graphs, we use text as a unifying medium. Many graphs in the real world contain textual features, such as citation networks and Wikipedia networks, and are known as Text-Attributed Graphs (TAGs). Compared to pre-processed vector features, textual features provide a consistent representation across different domains, which is beneficial for the transferability of the model. In addition to natural TAGs, we also explore using text to represent features of non-natural TAGs, such as molecular graphs, making it possible for the model to generalize to a broader range of domains.
@@
To learn transferable invariances across different graphs and tasks, we design a universal template to unify different tasks by contextualizing the nodes/edges/graphs on which we make predictions. Some perspectives suggest that representative local graph structural patterns are universal and transferable across different graphs (Newman, 2018). For node/edge-level tasks on large-scale graphs, we adopt the Personalized PageRank (PPR) algorithm to sample subgraphs, thereby mitigating differences in graph structures across different domains, while also making the model scalable.
@@
We propose a cascaded architecture of LMs and GNNs and new pre-training objectives based on Masked Graph Modeling (MGM), specifically tailored for self-supervised learning on TAGs at scale. To further endow our model with zero-shot capabilities, we concatenate pre-trained graph embeddings with natural language instructions and perform instruction tuning on LLMs, allowing us to unify the label spaces of graphs from different domains through natural language in a generative manner.
@@
  • We identify the challenges in developing a cross-domain graph foundation model and present the use of TAGs to unify graphs from diverse domains. We propose a foundation model for TAGs, named UniGraph, which incorporates a novel cascaded LM and GNN backbone. Additionally, we propose the first pre-training algorithm specifically designed for large-scale self-supervised learning on TAGs.
@@
  • We conduct experiments on node/edge/graph-level tasks across 11 different graph datasets from 5 distinct domains, with the largest graph comprising 111 million nodes. In cross-domain settings, UniGraph outperforms not just other cross-domain methods, but also supervised methods trained on the target dataset.
2. Related Work
Cross-graph learning. For graph transfer learning within the same domain, the most successful are pre-trained models on molecular graphs (Liu et al., 2021), benefiting from similar node/edge semantics. In addition, techniques such as fine-tuning (Hou et al., 2022; He et al., 2024b), domain adaptation (Dai et al., 2022), and prompt graphs (Huang et al., 2023) are used to achieve cross-graph transfer within the same domain. However, all these methods require the alignment of different graphs in both vector feature and label space. For cross-domain graph learning, GCC (Qiu et al., 2020) and ULTRA (Galkin et al., 2023), by ignoring node features and pretraining structural representations, achieve transfer learning for specific tasks, such as knowledge graph reasoning. OFA (Liu et al., 2023a) leverages pre-trained LMs to align the feature spaces of TAGs from different domains, while also transforming all downstream classification tasks into binary classification tasks, enabling it to conduct supervised training across all graphs.
Connections to existing methods. As a foundation model for TAGs, our UniGraph framework is distinct from OFA in three main aspects: 1) UniGraph performs end-to-end representation learning on TAGs, whereas OFA’s training is decoupled, using frozen LMs to generate features and then training GNNs separately. 2) UniGraph is a self-supervised learning framework, whereas OFA is a supervised learning framework that requires specific downstream task labels. 3) UniGraph, after pretraining, can generalize to any unseen target TAGs. In contrast, OFA co-trains on multiple target graphs simultaneously and then infers on each separately. Another type of existing graph learning models on TAGs are LLMs-only methods like LLaGA (Chen et al., 2024) and GraphGPT (Tang et al., 2023), which use instruction tuning to map graph data into the LLM embedding space. UniGraph can integrate with these methods, offering high-quality embeddings that improve LLMs’ understanding of graph structures. For conventional LMs+GNNs models that are not cross-domain models, they differ from UniGraph in both the target research problem and technical details. For example, GLEM (Zhao et al., 2022) and TAPE (He et al., 2024a) are supervised learning models on TAGs; they decouple LMs and GNNs, optimizing them separately. G2P2 (Wen and Fang, 2023) uses a CLIP (Radford et al., 2021)-like contrastive learning algorithm and employs prompt tuning to adapt to different downstream tasks.
Figure 1. Overview of UniGraph framework. 1) In pre-training, we employ a self-supervised approach, leveraging TAGs to unify diverse graph data. This phase involves a cascaded architecture combining LMs and GNNs. We propose Graph Siamese Masked Autoencoders as the training architecture, which learns to reconstruct the masked text of each node using the text of its neighbors. 2) In few-shot transfer, the pre-trained model can make predictions with minimal data by comparing the embeddings of the query and support graphs. 3) Zero-shot transfer is achieved through graph instruction tuning with LLMs, enabling it to understand category labels in natural language and make predictions on unseen graphs without any graph-specific training.
3. Preliminaries
3.1. Text-Attributed Graphs (TAGs)
@@
A Text-Attributed Graph (TAG) is defined as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T}_{\mathcal{V}},\mathcal{T}_{\mathcal{E}})$, where $\mathcal{V}$ represents the set of nodes and $\mathcal{E}$ represents the set of edges. For each node $v\in\mathcal{V}$, there is an associated text $t_{v}\in\mathcal{T}_{\mathcal{V}}$ representing node-level textual information. For each edge $e_{vu}\in\mathcal{E}$ connecting nodes $v$ and $u$, there is an associated text $t_{e_{vu}}\in\mathcal{T}_{\mathcal{E}}$ representing edge-level textual information. It is possible for a TAG to have only $\mathcal{T}_{\mathcal{V}}$.
3.2. Problem Definition
This paper introduces a cross-domain foundation model for TAGs, pre-trained in a self-supervised manner. Our model’s generalization ability is evaluated across domains using unseen datasets through three machine learning problems. The pre-trained model, denoted as $f_{\theta}$, operates on a TAG and generates an embedding for each node. Formally, the model’s function can be expressed as: $f_{\theta}:\mathcal{G}\rightarrow\mathbb{R}^{|V|\times d}$, where $d$ is the dimensionality of the embedding space.
Self-supervised representation learning. The primary aim of self-supervised learning is to produce embeddings that are useful in downstream tasks. We adopt a linear probing setting to evaluate the representation learning ability of a frozen pre-trained model $f_{\theta}$.
@@ 4. The UniGraph Approach
4.1. Unifying Graphs and Tasks from Diverse Domains
Graphs from different domains often have different applications, corresponding to different tasks. Graph learning tasks can generally be divided into node, edge, and graph-level tasks, each focusing on different parts of the graph. The key to using one model to handle any task on any graph lies in finding a universal function acting as a versatile mapping tool, adaptable to different graph learning tasks. In this paper, we utilize the concept of Anchor Node(s) and their contextual subgraph(s) to construct this universal function. The unification of node, edge, and graph-level tasks can be achieved through a general contextual subgraphs processing and Anchor Nodes embedding refinement operation, denoted by $\mathbf{h}=g(\mathcal{X},\mathcal{A})$, where $\mathbf{h}$ represents the output vector representation for a node, an edge or a graph, $\mathcal{X}$ denotes the set of contextual subgraphs, and $\mathcal{A}$ signifies the set of Anchor Node(s). Formally, we define the function $g$ with our pre-trained model $f_{\theta}$ using the format:
(1) $\mathbf{h}=g(\mathcal{X},\mathcal{A})=R(f_{\theta}(\mathcal{X}),\mathcal{A}),$
@@ 4.1. Unifying Graphs and Tasks from Diverse Domains
4.2. Graph Siamese Masked Autoencoders on TAGs
The existing graph self-supervised learning methods (Veličković et al., 2018b; Kipf and Welling, 2016b; Thakoor et al., 2021; Hou et al., 2022) adopt GNNs as the backbone networks and use pre-processed vector features as input. In this part, we propose an end-to-end self-supervised learning method on TAGs. We cascade a pre-trained language model (LM) and a GNN as the backbone network. In this paper, we adopt DeBERTa-base (He et al., 2020) as our LM and GAT (Veličković et al., 2018a) as our GNN. Inspired by successful self-supervised learning techniques such as masked modeling (Kenton and Toutanova, 2019; He et al., 2022) and siamese networks (Wu et al., 2018), we designed Graph Siamese Masked Autoencoders, enabling large-scale self-supervised pre-training on TAGs.
@@ 4.2. Graph Siamese Masked Autoencoders on TAGs
Encoder. Our encoder consists of an LM $f_{\text{LM}}$ and a GNN $f_{\text{GNN}}$. For each node $v\in\mathcal{V}$, the masked textual feature sequence $m_{v}$ is processed through the LM $f_{\text{LM}}$ to get the hidden representations: $\widetilde{{\bm{E}}}_{v}=f_{\text{LM}}(m_{v})$. Then, we extract embedding of the [CLS] token $\widetilde{{\bm{e}}}_{v_{\text{cls}}}\in\mathbb{R}^{d}$ from the masked output $\widetilde{{\bm{E}}}_{v}\in\mathbb{R}^{(n_{v}+2)\times d}$. The GNN $f_{\text{GNN}}$ propagates embeddings of the [CLS] tokens across the graph. The input to $f_{\text{GNN}}$ is a matrix $\widetilde{{\bm{E}}}_{\text{cls}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ consisting of embedding of the [CLS] token for all nodes. The output of the GNN, $\widetilde{{\bm{E}}}^{\prime}_{\text{cls}}\in\mathbb{R}^{|\mathcal{V}|\times d}$, is another matrix of embeddings, representing the propagated features: $\widetilde{{\bm{E}}}^{\prime}_{\text{cls}}=f_{\text{GNN}}(\mathcal{G},\widetilde{{\bm{E}}}_{\text{cls}})$
Decoder. In decoding, we adopt the masked language modeling (MLM) (Kenton and Toutanova, 2019) as objective. The intuition behind designing such training objectives is that the model can learn to reconstruct the masked text of each node using the text of its neighbors, thereby fully exploring the graph structure while learning to understand the text. For each node $v$, the embeddings of masked textual feature $\widetilde{{\bm{E}}}_{v}$ from the masked forward pass and the GNN output of [CLS] token $\widetilde{{\bm{e}}}^{\prime}_{v_{\text{cls}}}$ are concatenated and linearly transformed:
(3) ${\bm{H}}_{v}=\text{Linear}(\widetilde{{\bm{E}}}_{v}\oplus(\widetilde{{\bm{e}}}^{\prime}_{v_{\text{cls}}}\otimes\mathbf{1}_{n_{v}+2}^{\top})),$
@@
(4) $\mathcal{L}_{\text{mask}}=-\frac{1}{\sum_{v}\sum_{i}\mathbb{I}(v,i)}\sum_{v\in\mathcal{V}}\sum_{i=1}^{n_{v}}\mathbb{I}(v,i)\cdot\log\mathbf{P}_{v}[i,T_{i}]$
@@
Latent space regularization. Some recent works suggest that designing new pretext tasks in the latent space helps stabilize the training of masked autoencoders, further enhancing the quality of the latent space representations (Chen et al., 2023; Dong et al., 2022; Zhang et al., 2021b; He and Ma, 2022; Zhang et al., 2021a). In this work, we propose the use of a target network to impose regularization constraints on the latent space. The target network shares the same architecture as the encoder, comprising an LM $f^{\prime}_{\text{LM}}$ and a GNN $f^{\prime}_{\text{GNN}}$, with parameters not updated through gradients. The parameters of $f^{\prime}_{\text{LM}}(*;\delta^{\prime})$ and $f_{\text{LM}}(*;\delta)$ are shared, while the parameters of $f^{\prime}_{\text{GNN}}(*;\xi^{\prime})$ are updated through exponential moving average (EMA) of $f_{\text{GNN}}(*;\xi)$ using weight decay $\tau$ to avoid collapse in negative-free SSL frameworks (Grill et al., 2020; Caron et al., 2021): $\delta^{\prime}\leftarrow\delta\quad\text{and}\quad\xi^{\prime}\leftarrow\tau\xi^{\prime}+(1-\tau)\xi.$
@@
(5) $\mathcal{L}_{\text{latent}}=\frac{1}{|\mathcal{V}|}\sum_{i}^{|\mathcal{V}|}(1-\frac{\bar{{\bm{z}}}_{i}^{\top}{\bm{e}}^{\prime}_{{\text{cls}}_{i}}}{\lVert\bar{{\bm{z}}}_{i}\rVert\cdot\lVert{\bm{e}}^{\prime}_{{\text{cls}}_{i}}\rVert})$
@@
(6) $\mathcal{L}=\mathcal{L}_{\text{mask}}+\lambda\cdot\mathcal{L}_{\text{latent}}.$
@@
Table 1. Experiment results in self-supervised representation learning. We report accuracy (%) for node/edge classification tasks and ROC-AUC score (%) for graph classification tasks. The performance of supervised methods, which are trained on the individual target dataset, is marked in gray. UniGraph and other self-supervised baselines (rows in white) are pretrained on ogbn-Papers100M, and then evaluated on the individual target dataset. NoPretrain represents a randomly initialized model in our framework without any pre-training.